# Critical LAN and Score Tests for Mixed Fractional Models under High-Frequency Observation at $H=3/4$

### 24 Mar 2026 v1

@@ 4.2 CLT for the transformed score vector
$\left\lVert A_{n}^{\mathrm{m}}(u)\right\rVert_{\mathrm{op}}\leq\frac{|u_{1}|}{\sqrt{n\Delta_{n}L_{n}}}\left\lVert C_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}+\frac{|u_{2}|}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\left\lVert D_{n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}=O\!\left(\frac{1}{\sqrt{n\Delta_{n}L_{n}}}\right),$
whereas $\left\lVert A_{n}^{\mathrm{m}}(u)\right\rVert_{\mathrm{F}}^{2}\to 2u^{\top}\Gamma_{\mathrm{m}}^{\mathrm{crit}}u$. Since $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$ is positive definite, the operator/Frobenius ratio tends to 0 for every $u\neq 0$. By lemma 13,
@@
$\left\lVert A_{n}^{\mathrm{o}}(u)\right\rVert_{\mathrm{op}}\leq\frac{|u_{1}|}{\sqrt{n\Delta_{n}L_{n}}}\left\lVert C_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}+\frac{|u_{2}|}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\left\lVert D_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}+\frac{|u_{3}|}{\sqrt{n\Delta_{n}}}\left\lVert A_{n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}=O\!\left(\frac{1}{\sqrt{n\Delta_{n}}}\right).$
@@ 5.1 Matrix Taylor expansions
Diagonalize $S$. Since every eigenvalue belongs to [1/2,1/2][-1/2,1/2], the scalar Taylor expansions of $\log(1+x)$ and $(1+x)^{-1}$ can be summed over the spectrum term by term (see [3]). ∎
@@ 5.2 Critical local perturbation
$\left\lVert M_{\sigma\sigma,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{\sigma H,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}),$ $\left\lVert M_{HH,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}^{2}),$
@@
$(\delta_{\sigma,n}^{\mathrm{m}})^{2}\left\lVert M_{\sigma\sigma,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
$|\delta_{\sigma,n}^{\mathrm{m}}\delta_{H,n}^{\mathrm{m}}|\left\lVert M_{\sigma H,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
$(\delta_{H,n}^{\mathrm{m}})^{2}\left\lVert M_{HH,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),$
@@
$\left\lVert\mathcal{R}_{n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr),\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr)\to 0.$
@@
$\left\lVert M_{\sigma\sigma,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{\sigma H,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}),$ $\left\lVert M_{HH,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}^{2}),$
$\left\lVert M_{\sigma\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{H\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(L_{n}),$ $\left\lVert M_{\alpha\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$
@@
$(\delta_{\sigma,n}^{\mathrm{o}}\delta_{\alpha,n}^{\mathrm{o}})\left\lVert M_{\sigma\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}L_{n}^{-1/2}\bigr),$
$(\delta_{H,n}^{\mathrm{o}}\delta_{\alpha,n}^{\mathrm{o}})\left\lVert M_{H\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}L_{n}^{-1/2}\bigr),$
$(\delta_{\alpha,n}^{\mathrm{o}})^{2}\left\lVert M_{\alpha\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\bigl((n\Delta_{n})^{-1}\bigr),$
@@
$\left\lVert\mathcal{R}_{n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}=O\bigl((n\Delta_{n})^{-1}\bigr),\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)=O\bigl((n\Delta_{n})^{-1}\bigr)\to 0.$
Since equation 30 gives $\left\lVert B_{n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}\to 0$, the mfOU claim follows as well. ∎
@@ 5.3 Final proof of the LAN theorems
6 Score tests at the critical boundary
We now derive score-type tests from the critical LAN expansions established in theorems 9 and 10. Throughout, let
@@ 6 Score tests at the critical boundary
denote the $p$-quantile of the standard normal distribution.
Since the LAN expansions are centered at the boundary point $H=3/4$, the relevant tests are naturally calibrated under the critical constraint $H=3/4$. In the presence of nuisance parameters $\sigma$ and $\alpha$, the test for the $H$-direction is based on the efficient score, obtained by projecting the central sequence onto the orthocomplement of the nuisance tangent space.
(mfBm).
Let
@@
Treating $\sigma$ as nuisance, the efficient score for $H$ is
@@
and the corresponding efficient information is
@@
Accordingly, we consider the standardized statistic
@@
(mfOU).
Let
@@
be the normalized score vector in theorem 10. Write
$\eta:=(\sigma,\alpha),\qquad\Xi_{\eta,n}^{\mathrm{o}}:=\begin{pmatrix}\Xi_{\sigma,n}^{\mathrm{o}}\\[3.00003pt] \Xi_{\alpha,n}^{\mathrm{o}}\end{pmatrix}$,
and partition the critical information matrix as
$\Gamma_{\mathrm{o}}^{\mathrm{crit}}=\begin{pmatrix}\Gamma_{\eta\eta}^{\mathrm{o}}&\Gamma_{\eta H}^{\mathrm{o}}\\ \Gamma_{H\eta}^{\mathrm{o}}&\Gamma_{HH}^{\mathrm{o}}\end{pmatrix}$.
The efficient score for $H$ is then
@@
Since (11) yields $\Gamma_{H\alpha}^{\mathrm{o}}=0$, this reduces to
$\Delta_{H,n}^{\mathrm{eff,o}}=\Xi_{H,n}^{\mathrm{o}}-\frac{\sigma}{2}\Xi_{\sigma,n}^{\mathrm{o}}.$
Likewise,
$I_{H}^{\mathrm{eff,o}}:=\Gamma_{HH}^{\mathrm{o}}-\Gamma_{H\eta}^{\mathrm{o}}\bigl(\Gamma_{\eta\eta}^{\mathrm{o}}\bigr)^{-1}\Gamma_{\eta H}^{\mathrm{o}}=\frac{3\sigma^{4}}{64},$
and we define
@@
Assume the critical LAN expansions in theorems 9 and 10.
  1. (i)
    Under the critical null hypothesis $H=3/4$,
@@
  3. (iii)
    Therefore, an asymptotic level-$\alpha$ boundary calibrated one-sided test of
@@
    rejects for large values of the statistic:
@@
Under the critical null hypothesis $H=3/4$, the LAN central sequences in Theorems 9 and 10 satisfy
$\Xi_{n}^{\mathrm{m}}\xrightarrow[]{d}\mathcal{N}(0,\Gamma_{\mathrm{m}}^{\mathrm{crit}}),\qquad\Xi_{n}^{\mathrm{o}}\xrightarrow[]{d}\mathcal{N}(0,\Gamma_{\mathrm{o}}^{\mathrm{crit}}).$
We first treat the mfBm model. The nuisance parameter is $\sigma$. Hence the efficient $H$-score is the residual after projecting the $H$-component onto the nuisance direction
$\Delta_{H,n}^{\mathrm{eff,m}}:=\Xi^{\mathrm{m}}_{H,n}-\Gamma^{\mathrm{m}}_{H\sigma}\bigl(\Gamma^{\mathrm{m}}_{\sigma\sigma}\bigr)^{-1}\Xi^{\mathrm{m}}_{\sigma,n}.$
By the explicit entries of $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$,
$\Gamma^{\mathrm{m}}_{H\sigma}\bigl(\Gamma^{\mathrm{m}}_{\sigma\sigma}\bigr)^{-1}=\frac{\sigma}{2},$
so
@@
$I_{H}^{\mathrm{eff,m}}=\Gamma^{\mathrm{m}}_{HH}-\Gamma^{\mathrm{m}}_{H\sigma}\bigl(\Gamma^{\mathrm{m}}_{\sigma\sigma}\bigr)^{-1}\Gamma^{\mathrm{m}}_{\sigma H}.$
Therefore,
$T_{n}^{\mathrm{m}}=\frac{\Delta_{H,n}^{\mathrm{eff,m}}}{\sqrt{I_{H}^{\mathrm{eff,m}}}}\xrightarrow[]{d}\mathcal{N}(0,1).$
The mfOU case is analogous. Write
$\Xi_{n}^{\mathrm{o}}=\begin{pmatrix}\Xi^{\mathrm{o}}_{\eta,n}\\ \Xi^{\mathrm{o}}_{H,n}\end{pmatrix},\qquad\eta:=(\sigma,\alpha),\qquad\Gamma_{\mathrm{o}}^{\mathrm{crit}}=\begin{pmatrix}\Gamma^{\mathrm{o}}_{\eta\eta}&\Gamma^{\mathrm{o}}_{\eta H}\\ \Gamma^{\mathrm{o}}_{H\eta}&\Gamma^{\mathrm{o}}_{HH}\end{pmatrix}$.
Then
$\Delta_{H,n}^{\mathrm{eff,o}}:=\Xi^{\mathrm{o}}_{H,n}-\Gamma^{\mathrm{o}}_{H\eta}\bigl(\Gamma^{\mathrm{o}}_{\eta\eta}\bigr)^{-1}\Xi^{\mathrm{o}}_{\eta,n}.$
Since $\Gamma^{\mathrm{o}}_{H\alpha}=0$, this reduces to
@@
and its variance is
$I_{H}^{\mathrm{eff,o}}=\Gamma^{\mathrm{o}}_{HH}-\Gamma^{\mathrm{o}}_{H\eta}\bigl(\Gamma^{\mathrm{o}}_{\eta\eta}\bigr)^{-1}\Gamma^{\mathrm{o}}_{\eta H}.$
Hence
$T_{n}^{\mathrm{o}}=\frac{\Delta_{H,n}^{\mathrm{eff,o}}}{\sqrt{I_{H}^{\mathrm{eff,o}}}}\xrightarrow[]{d}\mathcal{N}(0,1).$
This proves part (i).
For part (ii), under the local alternatives in Theorems 9 and 10, the LAN expansion holds with central sequence $\Xi_{n}^{\mathrm{m}}$ or $\Xi_{n}^{\mathrm{o}}$. By Le Cam’s third lemma ([5]), under a local alternative with $H$-component $h_{H}$, the efficient score acquires the mean shift
$h_{H}I_{H}^{\mathrm{eff,m}}\qquad\text{or}\qquad h_{H}I_{H}^{\mathrm{eff,o}},$
while its variance remains $I_{H}^{\mathrm{eff,m}}$ or $I_{H}^{\mathrm{eff,o}}$, respectively. After standardization, this gives
Finally, part (iii) follows from part (i). Since the mean shift in part (ii) is positive when $h_{H}>0$, this is the natural right-tail rule for testing $H_{0}:H\leq 3/4$ against $H_{1}:H>3/4$. ∎
@@
Practical implementation.
The preceding corollaries are stated for the oracle boundary statistics $T_{n}^{\bullet}$, that is, the statistics obtained when $\sigma$, $\alpha$ are treated as known. For numerical simulatuon, For numerical simulation, the true nuisance parameters $\sigma$ and $\alpha$ are replaced by restricted estimators computed under the constraint $H=3/4$.
  1. 1.
    Fix $H=3/4$ and estimate the nuisance parameters under this restriction. For mfBm this yields a restricted estimator $\widehat{\sigma}_{0}$, and for mfOU one analogously obtains $(\widehat{\sigma}_{0},\widehat{\alpha}_{0})$.
  2. 2.
    Evaluate the score components at the restricted estimates and at $H=3/4$.
  3. 3.
    Form the normalized score vector and project it onto the efficient $H$-direction by the formulas above.
  4. 4.
    Standardize by the corresponding plug-in efficient information and compare the resulting statistic with the Gaussian critical value.
This plug-in construction is the numerical counterpart of the boundary-calibrated score test used in the simulation study. Its full asymptotic justification under a fully composite null hypothesis for all parameters is left for future work.
Remark 19.
For implementation, the oracle statistics above may be replaced by plug-in versions obtained after estimating the nuisance parameters under the restricted constraint $H=3/4$.
Remark 20.
At the boundary point H=3/4, the opposite one-sided direction is handled by the same statistic with a left-tail rejection region.
$T^{\bullet}_{n}<z_{\alpha},\qquad\bullet\in\{\mathrm{m},\mathrm{o}\}.$
7 Simulation study
This section presents a first finite sample study of the plug-in one-sided critical score test. We examine how the statistic behaves when the nuisance parameter is estimated under the restriction $H=3/4$, and whether it can distinguish the critical case $H=3/4$ from a supercritical alternative $H>3/4$. The aim is numerical illustration rather than a formal asymptotic justification under the composite null hypothesis.
We only simulate mfBm here. For the mfOU model, the critical score structure in the $(\sigma,H)$-block is the same, while the $\alpha$-direction is asymptotically orthogonal at the leading order (see (11)). Thus the main problem already appears in the mfBm model and how the restricted nuisance fit enters the feasible $H$-score test. We therefore report only mfBm simulations here and leave the mfOU case for future work.
In every replication we impose the null hypothesis restriction $H=3/4$ and compute the restricted Gaussian MLE

### 02 Apr 2026 v2

@@ 4.2 CLT for the transformed score vector
$\|A_{n}^{\mathrm{m}}(u)\|_{\mathrm{op}}\leq\frac{|u_{1}|}{\sqrt{n\Delta_{n}L_{n}}}\|C_{n}^{\mathrm{m}}\|_{\mathrm{op}}+\frac{|u_{2}|}{\sqrt{n\Delta_{n}}L_{n}^{3/2}}\|D_{n}^{\mathrm{m}}\|_{\mathrm{op}}=O\!\left(\frac{1}{\sqrt{n\Delta_{n}L_{n}}}\right)+O\!\left(\frac{\log n}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\right).$
Since
$\log n=L_{n}+\log(n\Delta_{n}),$
we have
$\frac{\log n}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}=O\!\left(\frac{1}{\sqrt{n\Delta_{n}L_{n}}}\right)+\frac{\log(n\Delta_{n})}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}=o(1),$
because $n\Delta_{n}\to\infty$ and $L_{n}\to\infty$. Therefore
$\|A_{n}^{\mathrm{m}}(u)\|_{\mathrm{op}}\to 0.$
In addition,
$\|A_{n}^{\mathrm{m}}(u)\|_{\mathrm{F}}^{2}\to 2u^{\top}\Gamma_{\mathrm{m}}^{\mathrm{crit}}u.$
Since $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$ is positive definite, the operator/Frobenius ratio tends to 0 for every $u\neq 0$. By lemma 13,
@@
$\|A_{n}^{\mathrm{o}}(u)\|_{\mathrm{op}}\leq\frac{|u_{1}|}{\sqrt{n\Delta_{n}L_{n}}}\|C_{n}^{\mathrm{o}}\|_{\mathrm{op}}+\frac{|u_{2}|}{\sqrt{n\Delta_{n}}L_{n}^{3/2}}\|D_{n}^{\mathrm{o}}\|_{\mathrm{op}}+\frac{|u_{3}|}{\sqrt{n\Delta_{n}}}\|A_{n}^{\mathrm{o}}\|_{\mathrm{op}}$
$O\!\left(\frac{1}{\sqrt{n\Delta_{n}L_{n}}}\right)+O\!\left(\frac{\log n}{\sqrt{n\Delta_{n}}\,L_{n}^{3/2}}\right)+O\!\left(\frac{1}{\sqrt{n\Delta_{n}}}\right)\xrightarrow[]{n\rightarrow\infty}0.$
@@ 5.1 Matrix Taylor expansions
Diagonalize $S$. Since every eigenvalue belongs to [1/2,1/2][-1/2,1/2], the scalar Taylor expansions of $\log(1+x)$ and $(1+x)^{-1}$ can be summed over the spectrum term by term (see Cai et al. (2026)). ∎
@@ 5.2 Critical local perturbation
$\left\lVert M_{\sigma\sigma,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{\sigma H,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O(\log n),$ $\left\lVert M_{HH,n}^{\mathrm{m}}\right\rVert_{\mathrm{op}}$ $=O((\log n)^{2}),$
@@
$(\delta_{\sigma,n}^{\mathrm{m}})^{2}\|M_{\sigma\sigma,n}^{\mathrm{m}}\|_{\mathrm{op}}$ $=O\!\left((n\Delta_{n}L_{n})^{-1}\right),$
$|\delta_{\sigma,n}^{\mathrm{m}}\delta_{H,n}^{\mathrm{m}}|\|M_{\sigma H,n}^{\mathrm{m}}\|_{\mathrm{op}}$ $=O\!\left(\frac{\log n}{n\Delta_{n}\,L_{n}^{2}}\right),$
$(\delta_{H,n}^{\mathrm{m}})^{2}\|M_{HH,n}^{\mathrm{m}}\|_{\mathrm{op}}$ $=O\!\left(\frac{(\log n)^{2}}{n\Delta_{n}\,L_{n}^{3}}\right).$
@@
$\left\lVert\mathcal{R}_{n}^{\mathrm{m}}(h)\right\rVert_{\mathrm{op}}=O\Bigl(\frac{(\log n)^{2}}{n\Delta_{n}\,L_{n}^{3}}\Bigr),\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{m}}(h)\bigr)^{2}\Bigr)=O\bigl((n\Delta_{n}L_{n})^{-1}\bigr)\to 0.$
@@
$\left\lVert M_{\sigma\sigma,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{\sigma H,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(\log n),$ $\left\lVert M_{HH,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O((\log n)^{2}),$
$\left\lVert M_{\sigma\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$ $\left\lVert M_{H\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(\log n),$ $\left\lVert M_{\alpha\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O(1),$
@@
$(\delta_{\sigma,n}^{\mathrm{o}}\delta_{\alpha,n}^{\mathrm{o}})\left\lVert M_{\sigma\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\Bigl(\frac{1}{n\Delta_{n}L^{1/2}_{n}}\Bigr),$
$(\delta_{H,n}^{\mathrm{o}}\delta_{\alpha,n}^{\mathrm{o}})\left\lVert M_{H\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\Bigl(\frac{\log n}{n\Delta_{n}L^{3/2}_{n}}\Bigr),$
$(\delta_{\alpha,n}^{\mathrm{o}})^{2}\left\lVert M_{\alpha\alpha,n}^{\mathrm{o}}\right\rVert_{\mathrm{op}}$ $=O\Bigl(\frac{1}{n\Delta_{n}}\Bigr),$
@@
$\left\lVert\mathcal{R}_{n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}=O\!\left(\frac{1+{(\log n)^{2}}/{L_{n}^{3}}}{n\Delta_{n}}\right),\qquad\mathrm{tr}\Bigl(\bigl(\mathcal{R}_{n}^{\mathrm{o}}(h)\bigr)^{2}\Bigr)=O\bigl((n\Delta_{n})^{-1}\bigr)\to 0.$
Since equation (29) gives $\left\lVert B_{n}^{\mathrm{o}}(h)\right\rVert_{\mathrm{op}}\to 0$, the mfOU claim follows as well.
@@ 5.3 Final proof of the LAN theorems
6 Boundary score tests and feasible implementation
The LAN expansions obtained above are useful only if they can be turned into statistics whose meaning is clear at the boundary. At $H=3/4$, this means more than writing down a standardized central sequence. The null hypothesis is composite on the left and nuisance estimation cannot be avoided in any serious application. What the critical geometry gives us is a particularly transparent solution to the first of these problems.
Once the score is expressed in the transformed coordinates of Section 2, the $H$-direction can be separated from the nuisance tangent space in exactly the same way as in regular Gaussian LAN problems, except that the relevant normalization now carries the logarithmic critical rate.
Throughout this section,
@@ 6 Boundary score tests and feasible implementation
denotes the $p$-quantile of the standard normal law. Because the local experiment is centered at the boundary itself, all critical values are calibrated at $H=3/4$. The econometric object of interest is therefore the $H$-specific part of the critical central sequence after removing the contribution of the nuisance parameter $\sigma$ in the mfBm model, and of the nuisance parameters $\sigma$ and $\alpha$ in the mfOU model.
In the mfBm experiment this construction is especially direct, because the only nuisance parameter is the scale $\sigma$. Let
@@
Projecting the $H$-component away from the nuisance direction yields the efficient central sequence
@@
with efficient information
@@
Accordingly, the natural boundary statistic is
@@
The formula is simple, but its interpretation matters. The statistic does not test whether the raw $H$-score is large in isolation. It tests whether the component of that score which remains after removing the critical scale effect points into the supercritical region.
The mfOU experiment is formally richer but conceptually no different. There the nuisance parameter is two-dimensional, with $\eta:=(\sigma,\alpha)$, and one might expect the additional drift direction to complicate the efficient score. The critical information matrix shows that this does not happen at the leading order. If
@@
denotes the normalized score vector in theorem 10, and if we write
$\Xi_{\eta,n}^{\mathrm{o}}:=\begin{pmatrix}\Xi_{\sigma,n}^{\mathrm{o}}\\[3.00003pt] \Xi_{\alpha,n}^{\mathrm{o}}\end{pmatrix},\qquad\Gamma_{\mathrm{o}}^{\mathrm{crit}}=\begin{pmatrix}\Gamma_{\eta\eta}^{\mathrm{o}}&\Gamma_{\eta H}^{\mathrm{o}}\\ \Gamma_{H\eta}^{\mathrm{o}}&\Gamma_{HH}^{\mathrm{o}}\end{pmatrix}$,
then the efficient $H$-score is
@@
Because (10) gives $\Gamma_{H\alpha}^{\mathrm{o}}=0$, the drift parameter falls out of the leading efficient projection and one obtains the same critical direction as in mfBm,
$\Delta_{H,n}^{\mathrm{eff,o}}=\Xi_{H,n}^{\mathrm{o}}-\frac{\sigma}{2}\Xi_{\sigma,n}^{\mathrm{o}},\qquad I_{H}^{\mathrm{eff,o}}=\frac{3\sigma^{4}}{64},$
so that
@@
This coincidence is not merely algebraic convenience. It shows that the critical inferential problem is driven by the same transformed $(\sigma,H)$ block in both models, while the additional OU drift enters only beyond the leading boundary geometry.
@@
Assume the critical LAN expansions in theorems 9 and 10. Under the critical null hypothesis $H=3/4$,
@@
Consequently, an asymptotic level-$\alpha$ boundary-calibrated one-sided test of
@@
rejects for large values of the statistic, that is, whenever
@@
Under the critical null, the LAN central sequences in theorems 9 and 10 converge to centered Gaussian limits with covariance matrices $\Gamma_{\mathrm{m}}^{\mathrm{crit}}$ and $\Gamma_{\mathrm{o}}^{\mathrm{crit}}$. In the mfBm experiment, efficient projection removes the scale direction from the $H$-component of the central sequence and leaves
@@
$I_{H}^{\mathrm{eff,m}}=\Gamma^{\mathrm{m}}_{HH}-\Gamma^{\mathrm{m}}_{H\sigma}\bigl(\Gamma^{\mathrm{m}}_{\sigma\sigma}\bigr)^{-1}\Gamma^{\mathrm{m}}_{\sigma H},$
so standardization yields $T_{n}^{\mathrm{m}}\xrightarrow[]{d}\mathcal{N}(0,1)$.
The mfOU case follows the same argument, except that the nuisance vector is $\eta=(\sigma,\alpha)$. Because the leading critical information satisfies $\Gamma^{\mathrm{o}}_{H\alpha}=0$, the efficient projection again reduces to
@@
and its variance is the corresponding Schur complement $I_{H}^{\mathrm{eff,o}}$. This gives $T_{n}^{\mathrm{o}}\xrightarrow[]{d}\mathcal{N}(0,1)$ under the boundary null.
Under local alternatives, Le Cam’s third lemma (van der Vaart (1998)) transports the LAN drift into the efficient score. A positive $H$-component therefore shifts the mean of the efficient statistic to the right while leaving the asymptotic variance unchanged, yielding the limits
$T_{n}^{m}\xrightarrow{d}N\!\bigl(h_{H}\sqrt{I_{H}^{\mathrm{eff},m}},1\bigr),\qquad T_{n}^{o}\xrightarrow{d}N\!\bigl(h_{H}\sqrt{I_{H}^{\mathrm{eff},o}},1\bigr).$
The one-sided rejection region follows immediately from this directional mean shift. ∎
The one-sided version is the economically relevant test throughout the paper, because the question of interest is whether the data have crossed into the supercritical region. A two-sided statistic can be written down just as easily, but it addresses a different question. We record it only for completeness.
@@
The oracle statistics $T_{n}^{\bullet}$, $\bullet\in\{\mathrm{m},\mathrm{o}\}$, obtained when $\sigma$ is known, are the right theoretical benchmarks, but they are not yet the objects that an applied researcher would compute. In practice one would estimate nuisance parameters under the boundary restriction $H=3/4$, evaluate the transformed score at that restricted estimator, and then standardize the efficient projection with the corresponding plug-in information. For mfBm this means replacing $\sigma$ by a restricted estimator $\widehat{\sigma}_{0}$, for mfOU one analogously works with $(\widehat{\sigma}_{0},\widehat{\alpha}_{0})$. The resulting statistic is the feasible boundary score.
This feasible version is where the composite null issue becomes genuinely econometric. The LAN calculations already show that the critical score itself has a clean nuisance orthogonal representation. What remains is to justify that estimating the nuisance parameters under the restriction $H=3/4$ does not feed back into the first-order null law. Put differently, one needs a result showing that the plug-in error enters only at a smaller order than the logarithmically normalized efficient score. In regular likelihood settings such a statement is often routine, here it is precisely the sort of point that deserves separate treatment because the critical rate is slow and the boundary experiment is singular before transformation.
Lemma 19.
Assume that under the boundary null $H=3/4$ the restricted mfBm estimator satisfies
$\widehat{\sigma}_{0}-\sigma=O_{P}\bigl((n\Delta_{n})^{-1/2}\bigr),$
and that there exists a deterministic sequence $\varepsilon_{n}\downarrow 0$ such that
$\sup_{|\widetilde{\sigma}-\sigma|\leq\varepsilon_{n}}\left|\partial_{\sigma}\widehat{T}_{n}^{\mathrm{m}}(\widetilde{\sigma})\right|=O_{P}\bigl((n\Delta_{n})^{1/2}L_{n}^{-3/2}\bigr).$
Then
$\widehat{T}_{n}^{\mathrm{m}}(\widehat{\sigma}_{0})=\widehat{T}_{n}^{\mathrm{m}}(\sigma)+o_{P}(1).$
In particular, once the oracle statistic has the standard normal null limit, the feasible mfBm statistic has the same first-order null law.
Proof.
With probability tending to one, $|\widehat{\sigma}_{0}-\sigma|\leq\varepsilon_{n}$, so the mean-value theorem yields
$\widehat{T}_{n}^{\mathrm{m}}(\widehat{\sigma}_{0})-\widehat{T}_{n}^{\mathrm{m}}(\sigma)=(\widehat{\sigma}_{0}-\sigma)\,\partial_{\sigma}\widehat{T}_{n}^{\mathrm{m}}(\widetilde{\sigma}_{n})$
for some random $\widetilde{\sigma}_{n}$ between $\widehat{\sigma}_{0}$ and $\sigma$. The first factor is $O_{P}((n\Delta_{n})^{-1/2})$ by assumption, whereas the second is uniformly $O_{P}((n\Delta_{n})^{1/2}L_{n}^{-3/2})$ on that neighborhood. Their product is therefore $O_{P}(L_{n}^{-3/2})=o_{P}(1)$, which proves the claim. ∎
This lemma gives the minimal condition needed to complete the feasible mfBm argument. Once the restricted estimator has the usual boundary rate and the feasible score is locally smooth in $\sigma$, the slow logarithmic normalization pushes the plug-in contribution below the first-order null scale. The mfOU case is handled in an analogous manner. There the nuisance vector is $(\sigma,\alpha)$ rather than $\sigma$ alone, but the same Taylor expansion applies componentwise and no new first-order term is created by the drift direction because the critical $(H,\alpha)$ block has already been projected out. This is why the Monte Carlo analysis below focuses on the feasible mfBm statistic rather than on the oracle object. The numerical question is not whether the LAN formulas can be standardized on paper, but whether the plug-in score already behaves like a practical diagnostic for entry into the supercritical regime.
The same construction can be turned around if one wishes to study movement below the boundary rather than above it. In that case the rejection region is placed in the left tail, that is, $T_{n}^{\bullet}<z_{\alpha}$ for $\bullet\in\{\mathrm{m},\mathrm{o}\}$.
7 Monte Carlo diagnostics for the feasible boundary statistic
The numerical question at the boundary is not whether the transformed score has the right sign on paper, but whether the feasible statistic can already be computed in a stable way and how conservative its null calibration remains once the logarithmic factor $L_{n}=\log(1/\Delta_{n})$ enters the normalization. We consider two mfBm experiments with 1000 replications each, both built around the feasible one-sided statistic.
where the nuisance parameter is estimated under the boundary restriction